OpenStudy (anonymous):

PLEASE HELP 16t^3/2t + 4 * t+2/8t

OpenStudy (anonymous):

\[\frac{ 16t^3 }{ 2t+4 } * \frac{ t+2 }{ 8t }\]

OpenStudy (anonymous):

factor and cancel \[\frac{16t^3(t+2)}{2(t+2)8t}\]

OpenStudy (anonymous):

or \[\frac{16t^3(t+2)}{16t(t+2)}\]

OpenStudy (anonymous):

what is left after you cancel?

OpenStudy (anonymous):

wait so you factor the numerators? like  ?

OpenStudy (anonymous):

no the numerator is already in factored form as \(8t(t+2)\)

OpenStudy (anonymous):

oh so after you cancel it's